OpenStudy (anonymous):

what is 18/20 minus 2/4 ?

OpenStudy (bahrom7893):

18/20 = 9/10 - 1/2 = .9-.5 = 0.4 = 2/5

OpenStudy (bahrom7893):

Please click on Become a fan if I helped, thanks!

OpenStudy (anonymous):

Oh man... I would not switch to decimals. Just find the common denominator of 10

OpenStudy (bahrom7893):

sorry i meant: 18/20 = 9/10 then: 9/10-1/2 = .9-.5=.4.. dunno easier for me, i just rewrote it in fraction just in case they needed it in fractions..

OpenStudy (anonymous):

right so just do 9/10 - 5/10 and you get 4/10. which reduces down to 2/5.

OpenStudy (bahrom7893):

oh well, lol everyone uses their methods..

OpenStudy (anonymous):

Remember, fractions are your friends!

OpenStudy (anonymous):

love fractions!